OpenStudy (anonymous):

i cant solve this problem in four years, aljon will be three times as old as he was 8 years ago. how old is he now?

OpenStudy (anonymous):

Suppose aljon age was x years, then x + 4 = 3 (x-8) Can you finish it?

OpenStudy (anonymous):

x+4 = 3 x -24 -2 x= -24-4 2x = 28 x=14

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

YW